OpenStudy (anonymous):

41. Cylinder A has a radius of 1 m and a height of 3 m. Cylinder B has a radius of 3 m and a height of 3 m. What is the ratio of the volume of cylinder A to the volume of cylinder B? (1 point) I think its 1:3

geerky42 (geerky42):

None, not 1:3 Do you know the volume of cylinder?

OpenStudy (anonymous):

Then 3:1? the answer choices are 1:1 1:3 1:9 3:1

geerky42 (geerky42):

You can start write ratio like this: \(\pi r^2h : \pi r^2h\)

geerky42 (geerky42):

which is same as \(r^2 h : r^2 h\), because \(\pi\)s cancel each other So what is radius of cylinder A? Height?

OpenStudy (anonymous):

so would it be 1:9? since they both go into 9

geerky42 (geerky42):

both? what do you mean?

geerky42 (geerky42):

How, I mean

OpenStudy (anonymous):

1 and 3, and 3 and 3

OpenStudy (anonymous):

1 and 3 is 3, 3 and 3 is 9

geerky42 (geerky42):

well, yes, but remember, radius is squared.

OpenStudy (anonymous):

so what would the answer be?

geerky42 (geerky42):

you will find this answer yourself, im here only to guide you... above, you plug in from cylinder A, for below, plug in from cylinder B